OpenStudy (anonymous):

Evaluate the function at the given values of the independent variable and simplify. g(x)=4x^2+3x-5.........g(x+2) and g(-x)

jimthompson5910 (jim_thompson5910):

g(x)=4x^2+3x-5 g(x+2)=4(x+2)^2+3(x+2)-5 g(x+2)=4(x^2+4x+4)+3(x+2)-5 g(x+2)=4x^2+16x+16+3x+6-5 g(x+2)=4x^2+19x+17 ------------ g(x)=4x^2+3x-5 g(-x)=4(-x)^2+3(-x)-5 g(-x)=4x^2+3(-x)-5 ..... Note: (-x)^2 = x^2 g(-x)=4x^2-3x-5